### v2

# Competent Discrete Time Modeling For analogue controlled PWM Converter Considering State-Feedback

Abstract
Ever since R.D.Middlebrook proposed the state space averaging notion. The small signal model has been widely used as a design tool to tune control parameters. As Moore’s law is continuing and the AI chip’s high demand for power consumption and dynamic response, the control bandwidth needs to be boosted. However, the average model has two basic assumptions: the low-frequency assumption, the small ripple assumption. In high-bandwidth design, these two assumptions are violated. In order to solve this, various methods have been proposed. This paper gives a comprehensive overview of the existing small signal model for PWM converters from the following perspectives: 1. model fidelity, 2. analytical tractability. 3. complexity of the derivation process and result 4.generality.
I Introduction
PWM converters are the fundamental part of power conversion since the American scholar William.E.Newell proposed the notion ”Power Electronics”. A Small Signal Linearized Model is required to design a quick and robust dynamic system. The state space averaging method is first proposed[1] [2]. The slide-averaging operator is introduced to simplify the model. However, two important assumptions are introduced:
  • Small-Ripple Assumption The switching ripple is well-attenuated in the PWM control system input. Thus, the PWM can be regarded as a pure gain term.
  • Low-frequency Assumption The control bandwidth is a lot lower than the switching frequency. Therefore, the side-band components in the system is neglected.
In the late 1970s and 1990s, the ripple-based control methods[3] [4] were introduced to enhance the dynamic response. Moreover, the high bandwidth Average-Mode Controlled converters are introduced. Therefore, two effects become significant:
  • Sideband (sampling) Effect: As shown on the figure, the sideband effect exists in the system, since the PWM modulator’s small signal behavior is equivalent to a sampler (dirac comb). As the bandwidth is boosted, the high-frequency components are injected. Then, the low-frequency assumption imposes significant restrictions.
  • Ripple Effect As less attenuation is introduced into the control system, the waveform in the comparator input has more side-band components than in the case of low-bandwidth average mode control. Then the low-ripple assumption imposes significant restrictions.
In the ripple-based controlled PWM converters, the two assumptions in the SSA methods are violated. In order to model the ripple-based control, various models are proposed. Either or both of these two effects are considered. In the left part of the Introduction, a brief review on the modelling method will be shown.
I-A Multi-frequency Method
I-A1 Two-frequency Model
The Fourier-Integral based describing function method is used in [5] to describe the frequency-coupling behavior. Therefore, the sideband-effect is considered. However, since only the lower-sideband ($fsw+fp$) generated by switching frequency and perturbation frequency is considered, it cannot explain the side-band effect competently. The ripple effect in the system is ignored. In terms of the analytical tractability, since the sidebands are introduced in the form of an iteration denominator, the model cannot show the analytical symbolic stability boundary. The derivation process is very complex.
I-A2 Four-fequency Model
This model is very similar to the Two-frequency Model. More sidebands are considered. And the ripple effect in modulation is considered. Because of more considered sidebands, it is more accurate. However, it is still not competent to fully explain the sideband effect. Moreover, as more sidebands are considered, the complexity of the derivation process of these types of Fourier-Integral-based methods increases significantly. This phenomena also happens on its result, which also further diminishes its analytical tractability compared to the Two-frequency Model.
I-A3 Matrix-Based Multi-Frequency Model (HSS)
The derivation process of the Fourier integral in [5] and [6] is very complex. However, the sideband effect requires more sidebands to be considered to enhance the model fidelity. Based on the sampling theorem (the Poisson summation formulae), the sideband mapping relation is derived easily without a complex Fourier integral. Furthermore, the derivation process is also simplified using the linear algebra tool. Therefore, the derivation process and the result of the model are expressed in a simpler form. Introducing arbitary numbers of sidebands become possible under this improvement. Moreover, the modeling result is extended to the MIMO form. The interaction (such as the beat frequency oscillation) between converters that are cascade connected can be explained by the model.
However, this model’s improvement only simplified the derivation process of the model. The model result is still in a complex iteration form. The analytical tractability of the model results is diminished even more.
I-B Sampled-data Method
I-C Closed-loop method
I-C1 Closed-loop discrete model with approximation
In [7], the digital small signal perturbation analysis is applied to the PCM-controlled BUCK converter. The voltage-second balancing approximation is introduced. In order to make the model more understandable, the pade approximation is introduced. Consequently, the model form is a simple Laplace-domain polynomial. So, the model form is very friendly to industry utilization. It is the first model that can help to predict sub-harmonic oscillations in a PCM-controlled converter effectively. The simplification using the Pade approximation enables the model to transfer from exponential(the Z operator) and polynomial hybrid form to a Laplace domain pure polynomial form. This class of models has gained wide acceptance among engineers, owing to its simple structure and Laplace-domain formulation, which align with conventional engineering practice. The derivation process of this class of models is often less rigorous than desired, and at times seems insufficiently motivated, especially for the introduction of ZOH. This limitation poses challenges for systematic extension and raises concerns regarding the strength of the underlying theoretical framework.
I-C2 Closed-loop Describing Function
This model [8] utilizes the same mathematical tools as the model in section I-A1. It focus on the time domain closed-loop relationship, therefore both the ripple and sideband effect are considered. In the derivation process, the author introduced the voltage-second balance approximation in [8] and ampere-second approximation in $V^{2}$ extension [9]. This category of models inherits the Pade approximation framework, thus preserving the the same practical advantages of simplicity and familiarity to engineers as the Closed-loop discrete model. In terms of the the forms of the result, the system is expressed as a polynomial in the Laplace domain. This representation also incorporates the concept of the linear two-port equivalent circuit, which was widely referenced during the era of averaged models, as a means of presenting the model results. Although this equivalent circuit provides a certain degree of physical intuition, it is in fact introduced by fitting the model output rather than being directly observed from the modeling process. The underlying derivation remains essentially a perturbation-based approach under closed-loop conditions. Unlike the Closed-loop discrete model, the derivation here is conducted with much greater rigor. This enhanced rigor makes it possible to extend the modeling framework without exposing the subsequent developments to the theoretical vulnerabilities of the Closed-loop discrete model. However, the derivation process of this type of model is still highly complex.
II Analytical Steady State modelling for converters using Poincaré map
We consider a piecewise linear time‐invariant system over one period $T_{s}$, divided into $n$ consecutive intervals of durations $T_{1},\dots,T_{n}$ (with $T_{s}=\sum_{i=1}^{n}T_{i}$). On interval $i$ the system is governed by
$\dot{\mathbf{X}}(t)=A_{i}\,\mathbf{X}(t)+B_{i}\,\mathbf{U},$
where the input $\mathbf{U}$ is held constant in each interval. Denote the state at the start of the $i$th interval by $\mathbf{X}_{i-1}$, then the discrete‐time update is
$\mathbf{X}_{i}=\Phi_{i}\,\mathbf{X}_{i-1}\;+\;\Gamma_{i},\quad\Phi_{i}=e^{A_{i}T_{i}},$ (1)
$\Gamma_{i}=\int_{0}^{T_{i}}e^{A_{i}(T_{i}-\tau)}\,B_{i}\,\mathbf{U}\,d\tau.$
If $A_{i}$ is invertible,
$\Gamma_{i}=\int_{0}^{T_{i}}e^{A_{i}(T_{i}-\tau)}\,B_{i}\,\mathbf{U}\,d\tau$ (2)
$=\int_{0}^{T}e^{A\tau}B_{i}U\,d\tau=A_{i}^{-1}\bigl(e^{A_{i}T}-I\bigr)B_{i}U.$
Our goal is to derive:
  1. 1.
    A closed‐form expression for $\mathbf{X}_{n}$ in terms of $\mathbf{X}_{0}$ and the $\Gamma_{i}$.
  2. 2.
    The fixed‐point equation for the periodic steady state $\mathbf{X}^{*}$ satisfying $\mathbf{X}_{n}=\mathbf{X}_{0}=\mathbf{X}^{*}$.
III Product Notation: Direction and Boundary
We introduce two notations for multiplying the transition matrices:
$\prod_{j=a}^{b}\!{}^{\rightarrow}\Phi_{j}\;:=\;\Phi_{a}\,\Phi_{a+1}\,\cdots\,\Phi_{b},\quad\prod_{j=a}^{b}\!{}^{\leftarrow}\Phi_{j}\;:=\;\Phi_{b}\,\Phi_{b-1}\,\cdots\,\Phi_{a},$
for $a\leq b$. In particular,
$\prod_{j=a}^{a}\!{}^{\rightarrow}\Phi_{j}=\prod_{j=a}^{a}\!{}^{\leftarrow}\Phi_{j}=\Phi_{a}.$
In our closed‐form formula only the “reverse” product $\prod_{j=1}^{n}{}^{\leftarrow}\Phi_{j}$ appears, avoiding any need for an empty‐product convention.
IV Main Results: Recursive Closed‐Form and Fixed‐Point Equation
IV-A Recursive Closed‐Form
For any $n\geq 1$, the state at the end of the $n$th interval is
$\boxed{\mathbf{X}_{n}=\Bigl(\prod_{j=1}^{n}{}^{\leftarrow}\Phi_{j}\Bigr)\,\mathbf{X}_{0}+\sum_{i=1}^{n-1}\Bigl(\prod_{j=i+1}^{n}{}^{\leftarrow}\Phi_{j}\Bigr)\,\Gamma_{i}+\Gamma_{n}}$
IV-B Periodic Fixed‐Point Equation
If a periodic steady state $\mathbf{X}^{*}=\mathbf{X}_{0}=\mathbf{X}_{n}$ exists, it satisfies
$\boxed{\bigl(I-\Pi\bigr)\,\mathbf{X}^{*}=\sum_{i=1}^{n-1}\Bigl(\prod_{j=i+1}^{\,n}{}^{\leftarrow}\Phi_{j}\Bigr)\,\Gamma_{i}+\Gamma_{n},}\quad\Pi:=\prod_{j=1}^{n}{}^{\leftarrow}\Phi_{j}.$
Take four state transition as an example:
$X_{1}$ $=\Phi_{1}X_{0}+\Gamma_{1},$ (2)
$X_{2}$ $=\Phi_{2}\Phi_{1}X_{0}+\Phi_{2}\,\Gamma_{1}+\Gamma_{2},$
$X_{3}$ $=\Phi_{3}\Phi_{2}\Phi_{1}X_{0}+\Phi_{3}\Phi_{2}\,\Gamma_{1}+\Phi_{3}\,\Gamma_{2}+\Gamma_{3},$
$X_{4}$ $=\Phi_{4}\Phi_{3}\Phi_{2}\Phi_{1}X_{0}+\Phi_{4}\Phi_{3}\Phi_{2}\,\Gamma_{1}+\Phi_{4}\Phi_{3}\,\Gamma_{2}+\Phi_{4}\,\Gamma_{3}+\Gamma_{4}.$
V Jacobian Linearization of PWL Switching Maps Under Four PWM Logics
This section formulates a unified Jacobian-based linearization for piecewise-linear (PWL) switching systems under four PWM logics: (i) constant-on-time (COT), (ii) constant-off-time (COFT), (iii) fixed-frequency trailing-edge (peak) PWM, and (iv) fixed-frequency leading-edge (valley) PWM. The differences among these cases are (a) the sampling phase (peak vs. valley), (b) the subinterval ordering in the Poincaré map, and (c) which subinterval durations are fixed or time-varying.
A. PWL explicit solution and timing derivatives
Consider a PWL segment $i$ governed by
$\dot{x}(t)=A_{i}x(t)+B_{i}U,$ (3)
where $x\in\mathbb{R}^{n}$ and $U$ is constant over the segment. The segment endpoint mapping over duration $T$ is
$x(T)$ $=\Phi_{i}(T)\,x(0)+\Gamma_{i}(T),$ (4)
$\Phi_{i}(T)$ $=e^{A_{i}T},$
$\Gamma_{i}(T)$ $=\int_{0}^{T}e^{A_{i}(T-\tau)}B_{i}U\,d\tau.$
If $A_{i}$ is invertible, $\Gamma_{i}(T)=A_{i}^{-1}\big(\Phi_{i}(T)-I\big)B_{i}U$ may be used; otherwise the integral definition in (4) remains valid.
Time derivatives required for Jacobian linearization are
$\frac{\partial\Phi_{i}(T)}{\partial T}$ $=A_{i}\Phi_{i}(T)=\Phi_{i}(T)A_{i},$ (5)
$\frac{\partial\Gamma_{i}(T)}{\partial T}$ $=\Phi_{i}(T)B_{i}U.$
These identities eliminate manual perturbation expansions of matrix exponentials.
B. Unified event (comparator) timing relation
Let the switching event be determined by a comparator quantity $y(t)=Kx(t)$ intersecting a ramp of slope $S_{e}$ against a control voltage $v_{c}$. At the switching edge,
$Kx_{\mathrm{edge}}=v_{c}+S_{e}\,\Delta T,$ (6)
where $\Delta T$ denotes the relevant timing variable (e.g., $T_{\mathrm{on}}$ or $T_{\mathrm{off}}$). Linearizing (6) yields
$\widehat{\Delta T}=\frac{K\hat{x}_{\mathrm{edge}}-\hat{v}_{c}}{S_{e}}.$ (7)
In the four cases below, the sampling instant is chosen at the switching edge so that $x_{\mathrm{edge}}$ coincides with the sampled state, enabling elimination of $\widehat{\Delta T}$ inside the linearized map.
V-A Constant-On-Time (COT): Valley-Triggered, Fixed $T_{\mathrm{on}}$
In COT, the on-time $T_{\mathrm{on}}$ is fixed while the off-time $T_{\mathrm{off},k}$ varies and is determined by the next valley-trigger event. Sampling is taken at the valley (turn-on) instant. Over one cycle,
$\text{on (fixed }T_{\mathrm{on}})\ \rightarrow\ \text{off (variable }T_{\mathrm{off},k}).$
A. Large-signal Poincaré map
Let $x_{k}$ denote the sampled state at the valley of cycle $k$. The one-step map is
$x_{k+1}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off},k})\Big(\Phi_{\mathrm{on}}(T_{\mathrm{on}})x_{k}+\Gamma_{\mathrm{on}}(T_{\mathrm{on}})\Big)+\Gamma_{\mathrm{off}}(T_{\mathrm{off},k}).$ (8)
B. Jacobian blocks
At the periodic steady state $(x^{\star},T_{\mathrm{off}}^{\star})$, define
$\Phi_{\mathrm{on}}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}),$ (9)
$\Phi_{\mathrm{off}}^{\star}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),$
$\Gamma_{\mathrm{on}}$ $=\Gamma_{\mathrm{on}}(T_{\mathrm{on}}),$
$\Gamma_{\mathrm{off}}^{\star}$ $=\Gamma_{\mathrm{off}}(T_{\mathrm{off}}^{\star}).$
The state Jacobian is
$\phi_{\mathrm{COT}}=\left.\frac{\partial x_{k+1}}{\partial x_{k}}\right|_{\star}=\Phi_{\mathrm{off}}^{\star}\,\Phi_{\mathrm{on}}.$ (10)
Using (5), the off-time sensitivity is
$\gamma_{\mathrm{off}}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off}}}\right|_{\star}$ (11)
$=A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}\Big(\Phi_{\mathrm{on}}x^{\star}+\Gamma_{\mathrm{on}}\Big)+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U.$
C. Eliminating $\widehat{T}_{\mathrm{off},k}$
The valley event occurs at the sampling instant, hence $x_{\mathrm{edge}}=x_{k+1}$ and $\Delta T=T_{\mathrm{off},k}$. Linearizing (8) gives
$\hat{x}_{k+1}=\phi_{\mathrm{COT}}\hat{x}_{k}+\Gamma_{\mathrm{off}}\widehat{T}_{\mathrm{off},k}.$ (12)
Substituting (7) yields
$\Big(I-\Gamma_{\mathrm{off}}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{COT}}\hat{x}_{k}-\Gamma_{\mathrm{off}}\hat{v}_{c,k+1}/S_{e}.$ (13)
V-B Constant-Off-Time (COFT): Peak-Triggered, Fixed $T_{\mathrm{off}}$
In COFT, the off-time $T_{\mathrm{off}}$ is fixed while the on-time $T_{\mathrm{on},k}$ varies and is determined by the next peak-trigger event. Sampling is taken at the peak (turn-off) instant. Over one cycle,
$\text{off (fixed }T_{\mathrm{off}})\ \rightarrow\ \text{on (variable }T_{\mathrm{on},k}).$
A. Large-signal Poincaré map
Let $x_{k}$ denote the sampled state at the peak of cycle $k$. The one-step map is
$x_{k+1}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on},k})\Big(\Phi_{\mathrm{off}}(T_{\mathrm{off}})x_{k}+\Gamma_{\mathrm{off}}(T_{\mathrm{off}})\Big)+\Gamma_{\mathrm{on}}(T_{\mathrm{on},k}).$ (14)
B. Jacobian blocks
At the periodic steady state $(x^{\star},T_{\mathrm{on}}^{\star})$, define
$\Phi_{\mathrm{off}}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}),$ (15)
$\Phi_{\mathrm{on}}^{\star}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),$
$\Gamma_{\mathrm{off}}$ $=\Gamma_{\mathrm{off}}(T_{\mathrm{off}}),$
$\Gamma_{\mathrm{on}}^{\star}$ $=\Gamma_{\mathrm{on}}(T_{\mathrm{on}}^{\star}).$
The state Jacobian is
$\phi_{\mathrm{COFT}}=\left.\frac{\partial x_{k+1}}{\partial x_{k}}\right|_{\star}=\Phi_{\mathrm{on}}^{\star}\,\Phi_{\mathrm{off}}.$ (16)
The on-time sensitivity is
$\gamma_{\mathrm{on}}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on}}}\right|_{\star}$ (17)
$=A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}\Big(\Phi_{\mathrm{off}}x^{\star}+\Gamma_{\mathrm{off}}\Big)+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U.$
C. Eliminating $\widehat{T}_{\mathrm{on},k}$
The peak event occurs at the sampling instant, hence $x_{\mathrm{edge}}=x_{k+1}$ and $\Delta T=T_{\mathrm{on},k}$. Linearizing (14) gives
$\hat{x}_{k+1}=\phi_{\mathrm{COFT}}\hat{x}_{k}+\Gamma_{\mathrm{on}}\widehat{T}_{\mathrm{on},k}.$ (18)
Substituting (7) yields
$\Big(I-\Gamma_{\mathrm{on}}K/S_{e}\Big)\hat{x}_{k+1}$ $=\phi_{\mathrm{COFT}}\hat{x}_{k}-\Gamma_{\mathrm{on}}\hat{v}_{c,k+1}/S_{e}.$ (19)
V-C Fixed-Frequency Trailing-Edge (Peak) PWM
In trailing-edge fixed-frequency PWM, the clock forces turn-on at the beginning of each switching period, while the comparator triggers turn-off (peak) within the period. The fixed-frequency constraint is
$T_{\mathrm{on},k}+T_{\mathrm{off},k}=T_{s}.$ (20)
To use the event equation at the sampling instant, sampling is chosen at the peak (turn-off) edge. A peak-to-peak map spans across the cycle boundary and depends on $T_{\mathrm{on},k}$ and $T_{\mathrm{on},k+1}$.
A. Large-signal peak-to-peak map
Let $x_{k}$ be sampled at the peak of cycle $k$. Then
$\text{off in cycle }k:\ T_{\mathrm{off},k}=T_{s}-T_{\mathrm{on},k},\quad\text{then on in cycle }k{+}1:\ T_{\mathrm{on},k+1}.$
Hence
$x_{k+1}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on},k+1})\Big(\Phi_{\mathrm{off}}(T_{s}-T_{\mathrm{on},k})x_{k}$ (21)
$\qquad\qquad+\Gamma_{\mathrm{off}}(T_{s}-T_{\mathrm{on},k})\Big)+\Gamma_{\mathrm{on}}(T_{\mathrm{on},k+1}).$
B. Jacobian blocks with two timing perturbations
At steady state $T_{\mathrm{on}}^{\star}$ and $T_{\mathrm{off}}^{\star}=T_{s}-T_{\mathrm{on}}^{\star}$, define
$\Phi_{\mathrm{on}}^{\star}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),$ (22)
$\Phi_{\mathrm{off}}^{\star}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),$
$\Gamma_{\mathrm{off}}^{\star}$ $=\Gamma_{\mathrm{off}}(T_{\mathrm{off}}^{\star}).$
The state Jacobian is
$\phi_{\mathrm{FF,pk}}=\left.\frac{\partial x_{k+1}}{\partial x_{k}}\right|_{\star}=\Phi_{\mathrm{on}}^{\star}\,\Phi_{\mathrm{off}}^{\star}.$ (23)
Define two timing sensitivities:
$\gamma_{+}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k+1}}\right|_{\star}$ (24)
$=A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}\Big(\Phi_{\mathrm{off}}^{\star}x^{\star}+\Gamma_{\mathrm{off}}^{\star}\Big)+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U,$
and, by the chain rule $\partial(T_{s}-T_{\mathrm{on},k})/\partial T_{\mathrm{on},k}=-1$,
$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k}}\right|_{\star}$ (25)
$=-\Phi_{\mathrm{on}}^{\star}\Big(A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}x^{\star}+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U\Big).$
Thus the linearized map is
$\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,pk}}\hat{x}_{k}+\gamma_{+}\widehat{T}_{\mathrm{on},k+1}+\gamma_{-}\widehat{T}_{\mathrm{on},k}.$ (26)
C. Eliminating $\widehat{T}_{\mathrm{on},k+1}$ at the peak edge
At the peak sampling instant, $\Delta T=T_{\mathrm{on},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$, hence
$\widehat{T}_{\mathrm{on},k+1}=\frac{K\hat{x}_{k+1}-\hat{v}_{c,k+1}}{S_{e}}.$ (27)
Substituting (27) into (26) yields
$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,pk}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{on},k}$ (28)
$\quad-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$
The appearance of $\widehat{T}_{\mathrm{on},k}$ is intrinsic to fixed-frequency peak sampling and maps to a $z^{-1}$ factor in the $z$-domain formulation.
V-D Fixed-Frequency Leading-Edge (Valley) PWM
In leading-edge fixed-frequency PWM, the clock forces turn-off at the beginning of each period, while the comparator triggers turn-on (valley) within the period. The fixed-frequency constraint (20) holds. Sampling is taken at the valley (turn-on) edge, yielding a valley-to-valley map that depends on $T_{\mathrm{off},k}$ and $T_{\mathrm{off},k+1}$.
A. Large-signal valley-to-valley map
Let $x_{k}$ be sampled at the valley of cycle $k$. Then
$\text{on in cycle }k:\ T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k},$
$\text{then off in cycle }k{+}1:\ T_{\mathrm{off},k+1}.$
Hence
$x_{k+1}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off},k+1})\Big(\Phi_{\mathrm{on}}(T_{s}-T_{\mathrm{off},k})x_{k}$ (29)
$\qquad\qquad+\Gamma_{\mathrm{on}}(T_{s}-T_{\mathrm{off},k})\Big)+\Gamma_{\mathrm{off}}(T_{\mathrm{off},k+1}).$
B. Jacobian blocks with two timing perturbations
At steady state $T_{\mathrm{off}}^{\star}$ and $T_{\mathrm{on}}^{\star}=T_{s}-T_{\mathrm{off}}^{\star}$, define
$\Phi_{\mathrm{on}}^{\star}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),$ (30)
$\Phi_{\mathrm{off}}^{\star}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),$
$\Gamma_{\mathrm{on}}^{\star}$ $=\Gamma_{\mathrm{on}}(T_{\mathrm{on}}^{\star}).$
The state Jacobian is
$\Phi_{\mathrm{FF,val}}=\left.\frac{\partial x_{k+1}}{\partial x_{k}}\right|_{\star}=\Phi_{\mathrm{off}}^{\star}\,\Phi_{\mathrm{on}}^{\star}.$ (31)
Define
$\gamma_{+}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k+1}}\right|_{\star}$ (32)
$=A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}\Big(\Phi_{\mathrm{on}}^{\star}x^{\star}+\Gamma_{\mathrm{on}}^{\star}\Big)+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U,$
and, using $\partial(T_{s}-T_{\mathrm{off},k})/\partial T_{\mathrm{off},k}=-1$,
$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k}}\right|_{\star}$ (33)
$=-\Phi_{\mathrm{off}}^{\star}\Big(A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}x^{\star}+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U\Big).$
Thus
$\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,val}}\hat{x}_{k}+\gamma_{+}\widehat{T}_{\mathrm{off},k+1}+\gamma_{-}\widehat{T}_{\mathrm{off},k}.$ (34)
C. Eliminating $\widehat{T}_{\mathrm{off},k+1}$ at the valley edge
At the valley sampling instant, $\Delta T=T_{\mathrm{off},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$, hence
$\widehat{T}_{\mathrm{off},k+1}=\frac{K\hat{x}_{k+1}-\hat{v}_{c,k+1}}{S_{e}}.$ (35)
Substituting (35) into (34) yields
$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,val}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{off},k}$ (36)
$\quad-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$
As in the trailing-edge case, the previous-cycle timing term $\widehat{T}_{\mathrm{off},k}$ maps to a $z^{-1}$ factor in the $z$-domain representation.
VI From Edge-Time Perturbation to Equivalent Duty Perturbation
[…]
A. Definition and normalization
Let $q(t)\in\{0,1\}$ denote the switch (or gating) function over a nominal clock grid with period $T_{s}$. The equivalent duty is the per-period normalized pulse area:
$d[n]\;=\;\frac{1}{T_{s}}\int_{nT_{s}}^{(n+1)T_{s}}q(t)\,dt,\qquad\hat{d}[n]\;=\;d[n]-d^{\star}.$ (37)
In frequency-domain manipulations, it is convenient to regard $\hat{d}(t)$ as a piecewise-constant (ZOH) reconstruction of the discrete sequence $\hat{d}[n]$; the final mapping below is therefore stated directly as an $s$-domain operator relating $\hat{d}(s)$ to the relevant timing perturbation.
B. Translation-type perturbation (COT/COFT): both edges drift together
In COT (fixed $T_{\mathrm{on}}$) or COFT (fixed $T_{\mathrm{off}}$), the event timing perturbation typically manifests as an accumulated drift of the pulse train relative to the nominal clock grid. Consider a pulse train whose $n$th pulse is translated by $\Delta t_{n}$ while its width $T_{w}$ is fixed (for COT, $T_{w}=T_{\mathrm{on}}$; for COFT, $T_{w}=T_{\mathrm{off}}$ depending on the chosen alignment edge):
$q(t)=\sum_{n\in\mathbb{Z}}\Big(H\!\big(t-(nT_{s}+\Delta t_{n})\big)$- (38)
$H\!\big(t-(nT_{s}+\Delta t_{n}+T_{w})\big)\Big),$
where $H(\cdot)$ is the Heaviside step. Linearizing w.r.t. the small shift $\Delta t_{n}$ (using $H(t-(a+\varepsilon))=H(t-a)-\varepsilon\,\delta(t-a)+o(\varepsilon)$) yields the distributional first-order variation
$\Delta q(t)$ $=q(t)-q^{\star}(t)$ (39)
$=\sum_{n\in\mathbb{Z}}\Delta t_{n}\Big(\delta\!\big(t-(nT_{s}+T_{w})\big)-\delta\!\big(t-nT_{s}\big)\Big).$
Taking Laplace transforms and performing the standard ZOH normalization (so that the resulting signal is an equivalent duty perturbation) gives the compact mapping
$\boxed{\frac{\hat{d}(s)}{\Delta t(s)}=-\frac{1}{T_{s}}\,\frac{1-e^{-sT_{w}}}{1-e^{-sT_{s}}}}$ (40)
where $\Delta t(s)$ denotes the Laplace transform of the (ZOH) reconstruction of the sequence $\Delta t_{n}$.
Low-frequency check.
As $s\to 0$, $(1-e^{-sT_{w}})/(1-e^{-sT_{s}})\to T_{w}/T_{s}$, hence
$\frac{\hat{d}(s)}{\Delta t(s)}\xrightarrow[s\to 0]{}-\frac{T_{w}}{T_{s}^{2}},$ (41)
which matches the intuitive static scaling “time shift / period” after accounting for the fact that a translation changes the overlap of each pulse with the fixed integration window $nT_{s},(n+1)T_{s})$.
C. Fixed-frequency PWM: only one edge moves (no future accumulation)
For fixed-frequency PWM, the clock enforces a strict period $T_{s}$ and locks one edge to the clock. Consequently, the duty perturbation depends only on the within-period movement of the free edge, and the mapping reduces to a simple scalar factor $1/T_{s}$ (up to a sign set by whether the pulse is lengthened or shortened).
1) Trailing-edge (peak) PWM.
The rising edge is clocked at $t=nT_{s}$, while the falling edge occurs at $t=nT_{s}+T_{\mathrm{on},n}$. A small falling-edge perturbation $\Delta t_{f,n}$ equivalently perturbs on-time by $\widehat{T}_{\mathrm{on},n}=\Delta t_{f,n}$, hence
$\boxed{\hat{d}[n]=\frac{\widehat{T}_{\mathrm{on},n}}{T_{s}}=\frac{\Delta t_{f,n}}{T_{s}}}\Longleftrightarrow\boxed{\hat{d}(s)=\frac{1}{T_{s}}\,\Delta t_{f}(s)}.$ (42)
2) Leading-edge (valley) PWM.
The falling edge is clocked at $t=nT_{s}$, and the rising edge occurs at $t=nT_{s}+T_{\mathrm{off},n}$; the switch then remains on until $t=(n+1)T_{s}$. A small rising-edge delay $\Delta t_{r,n}$ increases $T_{\mathrm{off},n}$ and therefore reduces duty:
$\boxed{\hat{d}[n]=-\frac{\widehat{T}_{\mathrm{off},n}}{T_{s}}=-\frac{\Delta t_{r,n}}{T_{s}}}\Longleftrightarrow\boxed{\hat{d}(s)=-\frac{1}{T_{s}}\,\Delta t_{r}(s)}.$ (43)
Remark (why fixed-frequency is “memoryless” in $\Delta t$).
In (40), the factor $(1-e^{-sT_{s}})^{-1}$ encodes the fact that a single timing perturbation shifts all future pulse start times via accumulation (typical of COT/COFT drift relative to a nominal grid). Under strict fixed-frequency operation, the clock resets the reference every period, and the duty depends only on the current-period edge displacement, yielding the scalar mappings in (42)–(43).
VII Li–Jian Distillation: Port-Structured Weak-Coupling Approximation and MIMO Balances
This section presents the Li–Jian distillation procedure, i.e., a matrix-based reduction that converts a fully coupled PWL state-space model into a port-structured MIMO form whose kernel is an integrator cascade. The reduced model (i) preserves the physically chosen output/port variable (through a known $C$-row), (ii) replaces internal resistive couplings by controlled-source cascades driven by ports, and (iii) exposes volt-second and amp-second balance as solvability / periodicity conditions of the distilled Poincaré map.
A. Original PWL segment model and the chosen physical port
Consider a two-state power stage (e.g., buck) with
$x(t)=\begin{bmatrix}i_{L}(t)\\ v_{C}(t)\end{bmatrix},\qquad\dot{x}(t)=A_{i}x(t)+B_{i,\mathrm{in}}\,v_{\mathrm{in}}(t),$ (44)
where $i$ indexes PWL subintervals (on/off, etc.), and $v_{\mathrm{in}}$ is piecewise constant inside each segment.
A key input to distillation is the chosen port/output scalar (measured or physically meaningful)
$v_{o}(t)=C_{\mathrm{phys}}\,x(t),$ (45)
where $C_{\mathrm{phys}}\in\mathbb{R}^{1\times 2}$ is known from the output network topology (e.g., load/cap-ESR divider). In other words, (45) declares what we mean by $v_{o}$” as a port variable.
B. Li–Jian rectification (state normalization) for an integrator cascade
The distillation is most transparent after a constant state transformation
$x_{r}=T_{r}x,\qquad A_{r,i}=T_{r}A_{i}T_{r}^{-1},\quad B_{r,i}=T_{r}B_{i,\mathrm{in}},\quad C_{r}=C_{\mathrm{phys}}T_{r}^{-1}.$ (46)
Here $T_{r}$ may be a sign-rectifier (e.g., $D_{r}=\mathrm{diag}(\pm 1,\pm 1)$) or a scaling/normalization. A commonly used choice in practice is the charge scaling
$T_{q}=\mathrm{diag}(1,C_{f}),\qquad x_{q}=\begin{bmatrix}i_{L}\\ q_{C}\end{bmatrix}=\begin{bmatrix}i_{L}\\ C_{f}v_{C}\end{bmatrix}$, (47)
because it makes the capacitor equation look like a pure integrator ($\dot{q}_{C}=i_{C}$). In this section we keep the physical state $x=[i_{L}\;v_{C}]^{\top}$ for clarity; the $q_{C}$ variant is obtained by (46)–(47).
C. Distilled kernel as an integrator cascade (what is being approximated)
The core modeling decision is to replace the segment-dependent “fully coupled” dynamics by a shared kernel that reflects ideal energy storage:
$A_{0}\;:=\;\begin{bmatrix}0&0\\ \frac{1}{C_{f}}&0\end{bmatrix}$. (48)
Equation (48) enforces two structural facts:
[…]
$\dot{v}_{C}$ $=\frac{1}{C_{f}}i_{L}\quad\text{(capacitor is an integrator of current)}.$
Thus, distillation does not “kill” coupling by deleting it; rather, it moves coupling out of $A_{i}$ into port-driven controlled-source terms so that the remaining kernel is an integrator cascade.
D. Port-structured MIMO form and the weak-coupling approximation
We now build a port-open MIMO segment model whose explicit inputs are
$u(t)=\begin{bmatrix}v_{\mathrm{in}}(t)\\ v_{o}(t)\end{bmatrix}$, (50)
and whose dynamics take the distilled form
$\dot{x}(t)=A_{0}x(t)+B_{u,i}\,v_{\mathrm{in}}(t)+B_{y,i}\,v_{o}(t).$ (51)
Here $B_{u,i},B_{y,i}\in\mathbb{R}^{2\times 1}$ are segment-dependent port injection vectors.
To connect (51) to the original coupled model (44), we also keep the physical port closure relation
$v_{o}=C_{\mathrm{phys}}x.$ (52)
Substituting (52) into (51) gives a rank-1 coupled closed form
$\dot{x}=\Big(A_{0}+B_{y,i}C_{\mathrm{phys}}\Big)x+B_{u,i}v_{\mathrm{in}}.$ (53)
Hence, distillation replaces a generic $2\times 2$ coupling matrix by a shared kernel $A_{0}$ plus a rank-1 port feedback $B_{y,i}C_{\mathrm{phys}}$.
Matrix extraction of $B_{y,i}$ (controlled-source identification).
Given $(A_{i},B_{i,\mathrm{in}},C_{\mathrm{phys}})$, one natural way to determine $B_{y,i}$ is: choose $B_{u,i}:=B_{i,\mathrm{in}}$ (exact input injection), and fit the remaining coupling by
$A_{i}-A_{0}\;\approx\;B_{y,i}C_{\mathrm{phys}}.$ (54)
Since $C_{\mathrm{phys}}$ is a row vector, (54) constrains $A_{i}-A_{0}$ to be approximated by a rank-1 matrix whose only “measurement” is $v_{o}=C_{\mathrm{phys}}x$. The least-squares (minimum Frobenius-norm residual) solution is obtained by right-multiplying by $C_{\mathrm{phys}}^{\top}$:
$B_{y,i}$ $=\Big(A_{i}-A_{0}\Big)C_{\mathrm{phys}}^{\top}\Big(C_{\mathrm{phys}}C_{\mathrm{phys}}^{\top}\Big)^{-1},$ (55)
$E_{i}$ $=\Big(A_{i}-A_{0}\Big)-B_{y,i}C_{\mathrm{phys}}.$
The weak-coupling approximation is precisely the modeling choice
$E_{i}\approx 0,$ (56)
i.e., all state-to-state coupling not representable through the chosen port $v_{o}$ is neglected. Engineering-wise, $E_{i}$ is small when parasitics (e.g., DCR/ESR) create only weak “hidden” couplings beyond what is visible at the chosen port.
Port-network interpretation (controlled-source cascades).
Equation (51) means:
  • The inductor equation is “voltage-source driven” by ports: $v_{\mathrm{in}}$ and $v_{o}$ appear as effective applied voltages.
  • The output network coupling is represented as a dependent source driven by $v_{o}$ (hence the rank-1 form $B_{y,i}C_{\mathrm{phys}}$ in (53)).
This is exactly the “partial controlled-source isolation” described: distillation pushes internal impedance effects into port-driven dependent sources, leaving $A_{0}$ as a clean integrator cascade.
E. MIMO Poincaré map under the distilled kernel (explicit, non-heuristic)
For each segment $i$ of duration $T_{i}$ under (51), define
$\Phi_{0}(T_{i}):=e^{A_{0}T_{i}},\qquad\Gamma_{i}(T_{i}):=\int_{0}^{T_{i}}e^{A_{0}(T_{i}-\tau)}\,\begin{bmatrix}B_{u,i}&B_{y,i}\end{bmatrix}d\tau,$ (57)
so that the segment endpoint map is
$x^{+}=\Phi_{0}(T_{i})x^{-}+\Gamma_{i}(T_{i})\,u,\qquad u=\begin{bmatrix}v_{\mathrm{in}}\\ v_{o}\end{bmatrix}$. (58)
Because $A_{0}$ in (48) is nilpotent with $A_{0}^{2}=0$, we have closed forms
$\Phi_{0}(T)$ $=I+A_{0}T,$ (59)
$\Gamma_{i}(T)$ $=\Big(TI+\tfrac{T^{2}}{2}A_{0}\Big)\begin{bmatrix}B_{u,i}&B_{y,i}\end{bmatrix}$.
Hence, the distilled map is explicit and fully matrix-algebraic.
F. How volt-second and amp-second balance appear in the MIMO self-consistency equation
Consider a two-segment cycle (buck on/off) with durations $T_{\mathrm{on}}$ and $T_{\mathrm{off}}$ (no duty ratio is introduced; timing is the primitive). The one-cycle map is obtained by composing (58):
$x_{k+1}$ $=\Phi_{0}(T_{\mathrm{off}})\Phi_{0}(T_{\mathrm{on}})x_{k}$ (60)
$\quad+\Phi_{0}(T_{\mathrm{off}})\Gamma_{\mathrm{on}}(T_{\mathrm{on}})\,u+\Gamma_{\mathrm{off}}(T_{\mathrm{off}})\,u$
$=:\Phi_{h}x_{k}+\Gamma_{h}u.$
A periodic steady state $x^{\star}$ satisfies the self-consistency equation
$\Big(I-\Phi_{h}\Big)x^{\star}=\Gamma_{h}u.$ (61)
(i) Volt-second balance as a solvability constraint (Fredholm-style, but EE-friendly).
For the kernel (48), using (59),
$\Phi_{h}=\Phi_{0}(T_{\mathrm{off}})\Phi_{0}(T_{\mathrm{on}})=\Big(I+A_{0}T_{\mathrm{off}}\Big)\Big(I+A_{0}T_{\mathrm{on}}\Big)=I+A_{0}T_{s},$ (62)
where $T_{s}=T_{\mathrm{on}}+T_{\mathrm{off}}$ and $A_{0}^{2}=0$. Thus
$I-\Phi_{h}=-A_{0}T_{s}=-\begin{bmatrix}0&0\\ \frac{T_{s}}{C_{f}}&0\end{bmatrix}$, (63)
which is singular. The first row of (61) becomes
$\underbrace{\begin{bmatrix}1&0\end{bmatrix}(I-\Phi_{h})}_{=\,0}\,x^{\star}=\begin{bmatrix}1&0\end{bmatrix}\Gamma_{h}u.$ (64)
Therefore, a periodic solution exists only if
$0$ $=\begin{bmatrix}1&0\end{bmatrix}\Gamma_{h}u$ (65)
$=\begin{bmatrix}1&0\end{bmatrix}\Big(\Phi_{0}(T_{\mathrm{off}})\Gamma_{\mathrm{on}}(T_{\mathrm{on}})+\Gamma_{\mathrm{off}}(T_{\mathrm{off}})\Big)u.$
Using (59) and the fact that $\begin{bmatrix}1&0\end{bmatrix}A_{0}=0$, one obtains the simple interpretation
$0=T_{\mathrm{on}}\underbrace{\begin{bmatrix}1&0\end{bmatrix}\begin{bmatrix}B_{u,\mathrm{on}}&B_{y,\mathrm{on}}\end{bmatrix}}_{\text{inductor-voltage gain}}u+$ (66)
$T_{\mathrm{off}}\underbrace{\begin{bmatrix}1&0\end{bmatrix}\begin{bmatrix}B_{u,\mathrm{off}}&B_{y,\mathrm{off}}\end{bmatrix}}_{\text{inductor-voltage gain}}u.$
Equation (66) is exactly the volt-second balance statement: the net “driving” of the inductor integrator over one cycle must be zero, otherwise $i_{L}$ cannot be periodic.
(ii) Amp-second balance as the equation that pins the periodic current level.
The second row of (61) reads, using (63),
$-\frac{T_{s}}{C_{f}}\,i_{L}^{\star}$ $=\begin{bmatrix}0&1\end{bmatrix}\Gamma_{h}u,$ (67)
which determines the DC level $i_{L}^{\star}$ required for the capacitor integrator to be periodic (net charge change over a cycle equals zero). This is the amp-second balance viewpoint: capacitor net current over a cycle must vanish, otherwise the capacitor-related integrator state drifts.
(iii) Why $v_{C}^{\star}$ can be free in the port-open MIMO form.
Since $I-\Phi_{h}$ in (63) has rank one, (61) cannot uniquely determine both components of $x^{\star}$: only the condition (65) and the current level in (67) are fixed, while $v_{C}^{\star}$ may remain free until the port closure (52) (or an external network / controller constraint) is imposed. This is precisely the $i_{L}$ has a solution but $v_{C}$ is free” behavior you observed.
G. Buck example (distilled MIMO directly shows both balances)
For an ideal buck with $v_{o}=v_{C}$ and a resistive load $R$ (represented as a port-dependent current sink in the capacitor equation), a distilled port-open segment model can be written as
$\dot{x}=A_{0}x+B_{u,i}v_{\mathrm{in}}+B_{y,i}v_{o},\qquad A_{0}=\begin{bmatrix}0&0\\ \frac{1}{C_{f}}&0\end{bmatrix}$, (68)
with
$\text{On:}\quad B_{u,\mathrm{on}}$ $=\begin{bmatrix}\frac{1}{L_{f}}\\[2.84526pt] 0\end{bmatrix}$, $B_{y,\mathrm{on}}$ $=\begin{bmatrix}-\frac{1}{L_{f}}\\[2.84526pt] -\frac{1}{RC_{f}}\end{bmatrix}$, (69)
$\text{Off:}\quad B_{u,\mathrm{off}}$ $=\begin{bmatrix}0\\[2.84526pt] 0\end{bmatrix}$, $B_{y,\mathrm{off}}$ $=\begin{bmatrix}-\frac{1}{L_{f}}\\[2.84526pt] -\frac{1}{RC_{f}}\end{bmatrix}$.
The first entries in (69) encode the inductor voltage: $v_{L}=v_{\mathrm{in}}-v_{o}$ (on) and $v_{L}=-v_{o}$ (off). The second entries encode the load current in the capacitor equation: $i_{C}=i_{L}-v_{o}/R$.
Applying (66) to (69) gives
$0$ $=T_{\mathrm{on}}\Big(\frac{1}{L_{f}}v_{\mathrm{in}}-\frac{1}{L_{f}}v_{o}\Big)+T_{\mathrm{off}}\Big(0-\frac{1}{L_{f}}v_{o}\Big)$ (70)
$=\frac{1}{L_{f}}\Big(T_{\mathrm{on}}v_{\mathrm{in}}-T_{s}v_{o}\Big),$
hence
$v_{o}=\frac{T_{\mathrm{on}}}{T_{s}}v_{\mathrm{in}}=D\,v_{\mathrm{in}},\qquad D:=\frac{T_{\mathrm{on}}}{T_{s}}.$ (71)
This is the classic buck volt-second result, obtained here as the existence condition of the periodic solution of the distilled MIMO map.
Next, the amp-second viewpoint (67) reduces (for this ideal case) to the capacitor net-current balance
$\overline{i_{C}}=\overline{i_{L}-\frac{v_{o}}{R}}=0\quad\Rightarrow\quad i_{L}^{\star}=\frac{v_{o}}{R},$ (72)
which pins the DC inductor current level consistent with a periodic capacitor state.
H. Summary (what distillation achieves)
Li–Jian distillation is a matrix-to-matrix reduction:
$(A_{i},B_{i,\mathrm{in}},C_{\mathrm{phys}})\quad\mapsto\quad(A_{0},B_{u,i},B_{y,i},C_{\mathrm{phys}}),$ (73)
where $A_{0}$ is a shared integrator-cascade kernel (48), and $B_{y,i}$ is extracted by the port projection (55). In the resulting port-open MIMO form (51), duty ratio does not appear: timing ($T_{\mathrm{on}},T_{\mathrm{off}}$) is the primitive. The classical volt-second and amp-second balances emerge automatically from the self-consistency equation (61) as (i) a solvability condition (65) and (ii) a periodic-current pinning equation (67).

### v3

@@
 V-C Fixed-Frequency Trailing-Edge (Peak) PWM
-In trailing-edge fixed-frequency PWM, the clock forces turn-on at the beginning of each switching period, while the comparator triggers turn-off (peak) within the period.
-The fixed-frequency constraint is
-$T_{\mathrm{on},k}+T_{\mathrm{off},k}=T_{s}.$ (20)
-To use the event equation at the sampling instant, sampling is chosen at the peak (turn-off) edge.
-A peak-to-peak map spans across the cycle boundary and depends on $T_{\mathrm{on},k}$ and $T_{\mathrm{on},k+1}$.
-A. Large-signal peak-to-peak map
-Let $x_{k}$ be sampled at the peak of cycle $k$. Then
-$\text{off in cycle }k:\ T_{\mathrm{off},k}=T_{s}-T_{\mathrm{on},k},\quad\text{then on in cycle }k{+}1:\ T_{\mathrm{on},k+1}.$
-Hence
-$x_{k+1}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on},k+1})\Big(\Phi_{\mathrm{off}}(T_{s}-T_{\mathrm{on},k})x_{k}$ (21)
-$\qquad\qquad+\Gamma_{\mathrm{off}}(T_{s}-T_{\mathrm{on},k})\Big)+\Gamma_{\mathrm{on}}(T_{\mathrm{on},k+1}).$
-B. Jacobian blocks with two timing perturbations
-At steady state $T_{\mathrm{on}}^{\star}$ and $T_{\mathrm{off}}^{\star}=T_{s}-T_{\mathrm{on}}^{\star}$, define
-$\Phi_{\mathrm{on}}^{\star}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),$ (22)
-$\Phi_{\mathrm{off}}^{\star}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),$
-$\Gamma_{\mathrm{off}}^{\star}$ $=\Gamma_{\mathrm{off}}(T_{\mathrm{off}}^{\star}).$
+In fixed-frequency trailing-edge PWM, the switch is turned on by the clock at the beginning of each cycle, while the comparator triggers turn-off within the cycle.
+The switching period is constant, hence for every cycle $k$,
+$T_{\mathrm{on},k}+T_{\mathrm{off},k}=T_{s},\qquad T_{\mathrm{off},k}=T_{s}-T_{\mathrm{on},k}.$ (20)
+Let the ramp have amplitude $V_{r}$ over one period $T_{s}$, so its slope is
+$S_{e}:=\frac{V_{r}}{T_{s}}.$ (21)
+V-C1 Large-signal peak-to-peak map
+Let $x_{k}$ be sampled at the peak (turn-off) edge of cycle $k$.
+From this sampling instant to the next peak, the trajectory consists of (i) the off-interval of cycle $k$ with duration $T_{\mathrm{off},k}$, followed by (ii) the on-interval of cycle k+1k{+}1 with duration $T_{\mathrm{on},k+1}$.
+Therefore,
+$x_{k+1}$ $=\Phi_{\mathrm{on}}\!\big(T_{\mathrm{on},k+1}\big)\Big(\Phi_{\mathrm{off}}\!\big(T_{\mathrm{off},k}\big)x_{k}+\Gamma_{\mathrm{off}}\!\big(T_{\mathrm{off},k}\big)\Big)$ (22)
+$\quad+\Gamma_{\mathrm{on}}\!\big(T_{\mathrm{on},k+1}\big),$
+where, for each mode $m\in\{\mathrm{on},\mathrm{off}\}$, $\Phi_{m}(T)=e^{A_{m}T}$ and $\Gamma_{m}(T)=\int_{0}^{T}e^{A_{m}(T-\tau)}B_{m}U\,d\tau$.
+V-C2 Jacobian blocks with two timing perturbations
+At the steady state, denote
+$T_{\mathrm{on}}^{\star},\quad T_{\mathrm{off}}^{\star}:=T_{s}-T_{\mathrm{on}}^{\star},\qquad x^{\star}.$ (23)
+Define
+$\Phi_{\mathrm{on}}^{\star}:=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),\quad\Phi_{\mathrm{off}}^{\star}:=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),\quad\Gamma_{\mathrm{off}}^{\star}:=\Gamma_{\mathrm{off}}(T_{\mathrm{off}}^{\star}).$ (24)
 The state Jacobian is
 $\Phi_{\mathrm{FF,pk}}=\left.\frac{\partial x_{k+1}}{\partial x_{k}}\right|_{\star}=\Phi_{\mathrm{on}}^{\star}\,\Phi_{\mathrm{off}}^{\star}.$ (25)
-Define two timing sensitivities:
+Define two timing sensitivities corresponding to the two durations that appear in (22). First, the sensitivity to the next-cycle on-time $T_{\mathrm{on},k+1}$:
 $\gamma_{+}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k+1}}\right|_{\star}$ (26)
-$=A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}\Big(\Phi_{\mathrm{off}}^{\star}x^{\star}+\Gamma_{\mathrm{off}}^{\star}\Big)+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U,$
-and, by the chain rule $\partial(T_{s}-T_{\mathrm{on},k})/\partial T_{\mathrm{on},k}=-1$,
-$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k}}\right|_{\star}$ (25)
+$=A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}\Big(\Phi_{\mathrm{off}}^{\star}x^{\star}+\Gamma_{\mathrm{off}}^{\star}\Big)+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U.$
+Second, the sensitivity to the current-cycle on-time $T_{\mathrm{on},k}$, which enters (22) only through $T_{\mathrm{off},k}=T_{s}-T_{\mathrm{on},k}$ in (20):
+$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k}}\right|_{\star}=-\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k}}\right|_{\star}$ (27)
 $=-\Phi_{\mathrm{on}}^{\star}\Big(A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}x^{\star}+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U\Big).$
@@
 $\hat{x}_{k+1}=\Phi_{\mathrm{FF,pk}}\hat{x}_{k}+\gamma_{+}\widehat{T}_{\mathrm{on},k+1}+\gamma_{-}\widehat{T}_{\mathrm{on},k}.$ (28)
-C. Eliminating $\widehat{T}_{\mathrm{on},k+1}$ at the peak edge
-At the peak sampling instant, $\Delta T=T_{\mathrm{on},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$, hence
+V-C3 Eliminating $\widehat{T}_{\mathrm{on},k+1}$ at the peak edge
+At the peak (turn-off) sampling instant, the edge time is $\Delta T=T_{\mathrm{on},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$. Using the standard ramp-comparator small-signal relation,
 $\widehat{T}_{\mathrm{on},k+1}=\frac{K\hat{x}_{k+1}-\hat{v}_{c,k+1}}{S_{e}}.$ (29)
 Substituting (29) into (28) yields
-$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,pk}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{on},k}$ (28)
-$\quad-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$
-The appearance of $\widehat{T}_{\mathrm{on},k}$ is intrinsic to fixed-frequency peak sampling and maps to a $z^{-1}$ factor in the $z$-domain formulation.
+$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,pk}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{on},k}-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$ (30)
 V-D Fixed-Frequency Leading-Edge (Valley) PWM
-In leading-edge fixed-frequency PWM, the clock forces turn-off at the beginning of each period, while the comparator triggers turn-on (valley) within the period.
-The fixed-frequency constraint (20) holds.
-Sampling is taken at the valley (turn-on) edge, yielding a valley-to-valley map that depends on $T_{\mathrm{off},k}$ and $T_{\mathrm{off},k+1}$.
-A. Large-signal valley-to-valley map
-Let $x_{k}$ be sampled at the valley of cycle $k$. Then
-$\text{on in cycle }k:\ T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k},$
-$\text{then off in cycle }k{+}1:\ T_{\mathrm{off},k+1}.$
-Hence
-$x_{k+1}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off},k+1})\Big(\Phi_{\mathrm{on}}(T_{s}-T_{\mathrm{off},k})x_{k}$ (29)
-$\qquad\qquad+\Gamma_{\mathrm{on}}(T_{s}-T_{\mathrm{off},k})\Big)+\Gamma_{\mathrm{off}}(T_{\mathrm{off},k+1}).$
-B. Jacobian blocks with two timing perturbations
-At steady state $T_{\mathrm{off}}^{\star}$ and $T_{\mathrm{on}}^{\star}=T_{s}-T_{\mathrm{off}}^{\star}$, define
-$\Phi_{\mathrm{on}}^{\star}$ $=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),$ (30)
-$\Phi_{\mathrm{off}}^{\star}$ $=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),$
-$\Gamma_{\mathrm{on}}^{\star}$ $=\Gamma_{\mathrm{on}}(T_{\mathrm{on}}^{\star}).$
+In fixed-frequency leading-edge PWM, the clock forces turn-off at the beginning of each cycle, while the comparator triggers turn-on (valley) within the cycle.
+The fixed-period constraint implies, for every cycle $k$,
+$T_{\mathrm{on},k}+T_{\mathrm{off},k}=T_{s},\qquad T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}.$ (31)
+The ramp slope $S_{e}$ is still given by (21).
+V-D1 Large-signal valley-to-valley map
+Let $x_{k}$ be sampled at the valley (turn-on) edge of cycle $k$.
+From this sampling instant to the next valley, the trajectory consists of (i) the on-interval of cycle $k$ with duration $T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}$, followed by (ii) the off-interval of cycle k+1k{+}1 with duration $T_{\mathrm{off},k+1}$.
+Therefore,
+$x_{k+1}$ $=\Phi_{\mathrm{off}}\!\big(T_{\mathrm{off},k+1}\big)\Big(\Phi_{\mathrm{on}}\!\big(T_{\mathrm{on},k}\big)x_{k}+\Gamma_{\mathrm{on}}\!\big(T_{\mathrm{on},k}\big)\Big)$ (32)
+$\quad+\Gamma_{\mathrm{off}}\!\big(T_{\mathrm{off},k+1}\big),$
+with $T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}$ per (31).
+V-D2 Jacobian blocks with two timing perturbations
+At the steady state, denote
+$T_{\mathrm{off}}^{\star},\quad T_{\mathrm{on}}^{\star}:=T_{s}-T_{\mathrm{off}}^{\star},\qquad x^{\star}.$ (33)
+Define
+$\Phi_{\mathrm{on}}^{\star}:=\Phi_{\mathrm{on}}(T_{\mathrm{on}}^{\star}),\quad\Phi_{\mathrm{off}}^{\star}:=\Phi_{\mathrm{off}}(T_{\mathrm{off}}^{\star}),\quad\Gamma_{\mathrm{on}}^{\star}:=\Gamma_{\mathrm{on}}(T_{\mathrm{on}}^{\star}).$ (34)
 The state Jacobian is
 $\Phi_{\mathrm{FF,val}}=\left.\frac{\partial x_{k+1}}{\partial x_{k}}\right|_{\star}=\Phi_{\mathrm{off}}^{\star}\,\Phi_{\mathrm{on}}^{\star}.$ (35)
-Define
+Define the two timing sensitivities. First, the sensitivity to the next-cycle off-time $T_{\mathrm{off},k+1}$:
 $\gamma_{+}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k+1}}\right|_{\star}$ (36)
-$=A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}\Big(\Phi_{\mathrm{on}}^{\star}x^{\star}+\Gamma_{\mathrm{on}}^{\star}\Big)+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U,$
-and, using $\partial(T_{s}-T_{\mathrm{off},k})/\partial T_{\mathrm{off},k}=-1$,
-$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k}}\right|_{\star}$ (33)
+$=A_{\mathrm{off}}\Phi_{\mathrm{off}}^{\star}\Big(\Phi_{\mathrm{on}}^{\star}x^{\star}+\Gamma_{\mathrm{on}}^{\star}\Big)+\Phi_{\mathrm{off}}^{\star}B_{\mathrm{off}}U.$
+Second, the sensitivity to the current-cycle off-time $T_{\mathrm{off},k}$ which enters (32) only through $T_{\mathrm{on},k}=T_{s}-T_{\mathrm{off},k}$ in (31):
+$\gamma_{-}$ $=\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{off},k}}\right|_{\star}=-\left.\frac{\partial x_{k+1}}{\partial T_{\mathrm{on},k}}\right|_{\star}$ (37)
 $=-\Phi_{\mathrm{off}}^{\star}\Big(A_{\mathrm{on}}\Phi_{\mathrm{on}}^{\star}x^{\star}+\Phi_{\mathrm{on}}^{\star}B_{\mathrm{on}}U\Big).$
@@
 $\hat{x}_{k+1}=\Phi_{\mathrm{FF,val}}\hat{x}_{k}+\gamma_{+}\widehat{T}_{\mathrm{off},k+1}+\gamma_{-}\widehat{T}_{\mathrm{off},k}.$ (38)
-C. Eliminating $\widehat{T}_{\mathrm{off},k+1}$ at the valley edge
-At the valley sampling instant, $\Delta T=T_{\mathrm{off},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$, hence
+V-D3 Eliminating $\widehat{T}_{\mathrm{off},k+1}$ at the valley edge
+At the valley (turn-on) sampling instant, the edge time is $\Delta T=T_{\mathrm{off},k+1}$ and $x_{\mathrm{edge}}=x_{k+1}$. Thus
 $\widehat{T}_{\mathrm{off},k+1}=\frac{K\hat{x}_{k+1}-\hat{v}_{c,k+1}}{S_{e}}.$ (39)
 Substituting (39) into (38) yields
-$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,val}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{off},k}$ (36)
-$\quad-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$
-As in the trailing-edge case, the previous-cycle timing term $\widehat{T}_{\mathrm{off},k}$ maps to a $z^{-1}$ factor in the $z$-domain representation.
+$\Big(I-\gamma_{+}K/S_{e}\Big)\hat{x}_{k+1}$ $=\Phi_{\mathrm{FF,val}}\hat{x}_{k}+\gamma_{-}\widehat{T}_{\mathrm{off},k}-\gamma_{+}\hat{v}_{c,k+1}/S_{e}.$ (40)
 VI From Edge-Time Perturbation to Equivalent Duty Perturbation
@@
 Under strict fixed-frequency operation, the clock resets the reference every period, and the duty depends only on the current-period edge displacement, yielding the scalar mappings in (46)–(47).
-VII Li–Jian Distillation: Port-Structured Weak-Coupling Approximation and MIMO Balances
-This section presents the Li–Jian distillation procedure, i.e., a matrix-based reduction that converts a fully coupled PWL state-space model into a port-structured MIMO form whose kernel is an integrator cascade.
+VII Jian–Li Distillation: Port-Structured Weak-Coupling Approximation and MIMO Balances
+This section presents the Jian–Li distillation procedure, i.e., a matrix-based reduction that converts a fully coupled PWL state-space model into a port-structured MIMO form whose kernel is an integrator cascade.
 The reduced model (i) preserves the physically chosen output/port variable (through a known $C$-row), (ii) replaces internal resistive couplings by controlled-source cascades driven by ports, and (iii) exposes volt-second and amp-second balance as solvability / periodicity conditions of the distilled Poincaré map.